OpenStudy (anonymous):

are theses right????? What is the quotient: (2x2 + 6x – 8) ÷ (x + 4) ? 2x – 6 2x + 6 2x – 2 --->MY ANSWER 2x + 2 According to The Fundamental Theorem of Algebra, how many zeros does the function f(x) = 4x5 – 2x – 15 have? 2 3 4 5 --->MY ANSWER

OpenStudy (phi):

look good

OpenStudy (anonymous):

thanks for checking

OpenStudy (phi):

\[ \frac{2x^2 + 6x – 8}{x+4} = \frac{2(x-1)(x+4)}{(x+4) }= 2(x-1)= 2x-2\]

OpenStudy (phi):

for the 2nd question, the highest order term (biggest exponent) is 4x^5 that means you have 5 "roots" (though they may be complex or repeated )

OpenStudy (anonymous):

thanks :)